### v1

# Non-isotropic translation and modulation invariant Hilbert spaces

Abstract.
Let $\mathcal{H}$ be a Hilbert space of distributions on $\mathbf{R}^{d}$ which contains at least one non-zero element in $\mathscr{D}^{\prime}(\mathbf{R}^{d})$. If there is a constant $C_{0}>0$ such that
$\|e^{i\langle\,\cdot\,,\xi\rangle}f(\,\cdot\,-x)\|_{\mathcal{H}}\leq C_{0}\|f\|_{\mathcal{H}},\qquad f\in\mathcal{H},\ x,\xi\in\mathbf{R}^{d},$
then we prove that $\mathcal{H}=L^{2}(\mathbf{R}^{d})$, with equivalent norms.
Key words and phrases:
modulation spaces, Feichtinger’s minimization principle
1991 Mathematics Subject Classification:
46C15, 46C05, 42B35
0. Introduction
In [15] it is proved that a suitable non-trivial Hilbert space $\mathcal{H}\subseteq\mathscr{S}^{\prime}(\mathbf{R}^{d})$, which is norm preserved under translations
$\big{(}\,y\mapsto f(y)\,\big{)}$ $\mapsto\big{(}\,y\mapsto f(y-x)\,\big{)}$
and modulations
$\big{(}\,y\mapsto f(y)\,\big{)}$ $\mapsto\big{(}\,y\mapsto e^{i\langle y,\xi\rangle}f(y)\,\big{)}$
is equal to $L^{2}(\mathbf{R}^{d})$. (See [9] or Section 1 for notations.) It is here also proved that the norms between $\mathcal{H}$ and $L^{2}(\mathbf{R}^{d})$ only differs by a multiplicative constant, i. e. for some constant $C>0$ one has
$\|f\|_{\mathcal{H}}=C\|f\|_{L^{2}(\mathbf{R}^{d})},\qquad f\in\mathcal{H}=L^{2}(\mathbf{R}^{d}).$ (0.1)
This property is analogous to Feichtinger’s minimization property, which shows that the Feichtinger algebra $S_{0}(\mathbf{R}^{d})$, which is the same as the modulation space $M^{1,1}(\mathbf{R}^{d})$, is the smallest non-trivial Banach space of tempered distributions which is norm invariant under translations and modulations. (See e. g. [7] for general facts about modulation spaces.)
The condition on non-triviality in [15] is simply that $\mathcal{H}$ should contain at least one non-trivial element in $M^{1,1}(\mathbf{R}^{d})$. Hence the main result in [15] can be formulated as follows.
Theorem 0.1.
Let $\mathcal{H}$ be a Hilbert space which is continuously embedded in $\mathscr{S}^{\prime}(\mathbf{R}^{d})$, and norm preserved under translations and modulations. If $\mathcal{H}$ contains a non-zero element in $M^{1,1}(\mathbf{R}^{d})$, then $\mathcal{H}=L^{2}(\mathbf{R}^{d})$, and (0.1) holds for some constant $C>0$ which is independent of $f\in\mathcal{H}=L^{2}(\mathbf{R}^{d})$.
An alternative approach, using the Bargmann transform, to reach similar properties is given by Bais, Pinlodi and Venku Naidu in [1]. In fact, by using their result [1, Theorem 3.1] in combination with some well-known arguments in the distribution theory, one obtains the following improvement of Theorem 0.1.
Theorem 0.2.
Let $\mathcal{H}$ be a Hilbert space which is continuously embedded in $\mathscr{S}^{\prime}(\mathbf{R}^{d})$, and norm preserved under translations and modulations. If $\mathcal{H}$ is non-trivial, then $\mathcal{H}=L^{2}(\mathbf{R}^{d})$, and (0.1) holds for some constant $C>0$ which is independent of $f\in\mathcal{H}=L^{2}(\mathbf{R}^{d})$.
We observe that the condition
$|(f,\phi)|\leq C\|f\|_{\mathcal{H}},\qquad f\in\mathcal{H},\ \phi(x)=e^{-\frac{1}{2}|x|^{2}},$ (0.2)
for some constant $C>0$ in the hypothesis in [1, Theorem 3.1] is absent in Theorem 0.2. Hence Theorem 0.2 is slightly more general than [1, Theorem 3.1].
Here we remark that we may relax the hypothesis in Theorem 0.2 by assuming that $\mathcal{H}$ is continuously embedded in the space $\mathscr{D}^{\prime}(\mathbf{R}^{d})$ instead of the smaller space $\mathscr{S}^{\prime}(\mathbf{R}^{d})$. In fact, let $\mathcal{B}$ be a translation invariant Banach space which is continuously embedded in $\mathscr{D}^{\prime}(\mathbf{R}^{d})$ and satisfies
$\|f(\,\cdot\,-x)\|_{\mathcal{B}}\leq C\|f\|_{\mathcal{B}},$
for some constant $C\geq 1$ which is independent of $f\in\mathcal{B}$ and $x\in\mathbf{R}^{d}$. Then it follows by some standard arguments that $\mathcal{B}$ is continuously embedded in $\mathscr{S}^{\prime}(\mathbf{R}^{d})$ (see e. g. [15, Proposition 1.5]).
In the paper we investigate properties on weaker forms of translation and modulation invariant Hilbert spaces compared to [1, 15]. More precisely we show that except for the norm identity (0.1), Theorem 0.2 still holds true after the condition that $\mathcal{H}$ is norm preserved under translations and modulations, is relaxed into the weaker condition
$\|f(\,\cdot\,-x)e^{i\langle\,\cdot\,,\xi\rangle}\|_{\mathcal{H}}\leq C_{0}\|f\|_{\mathcal{H}},\qquad f\in\mathcal{H},\ x,\xi\in\mathbf{R}^{d}.$ (0.3)
In our extension, the identity (0.1) should be replaced by the norm equivalence
$C_{0}^{-1}C\|f\|_{L^{2}(\mathbf{R}^{d})}\leq\|f\|_{\mathcal{H}}\leq C_{0}C\|f\|_{L^{2}(\mathbf{R}^{d})},\qquad f\in\mathcal{H}=L^{2}(\mathbf{R}^{d}).$ (0.4)
More precisely our extension of Theorem 0.2 is the following.
Theorem 0.3.
Let $\mathcal{H}$ be a Hilbert space which is continuously embedded in $\mathscr{S}^{\prime}(\mathbf{R}^{d})$, and such that (0.3) holds true for some constant $C_{0}\geq 1$ which is independent of $f\in\mathcal{H}$ and $x,\xi\in\mathbf{R}^{d}$. If $\mathcal{H}$ is non-trivial, then $\mathcal{H}=L^{2}(\mathbf{R}^{d})$, and (0.4) holds for some constant $C>0$ which is independent of $f\in\mathcal{H}=L^{2}(\mathbf{R}^{d})$.
In Section 2 we also present some improvements of this result, where it is assumed that $\mathcal{H}$ in Theorem 0.3 is embedded in suitable larger (ultra-)distribution spaces which contain $\mathscr{S}^{\prime}(\mathbf{R}^{d})$ (see Theorem 2.1). These investigations are based on some general properties of translation and modulation invariant quasi-Banach spaces (see Proposition 2.2).
Acknowledgement
The authors are grateful to HRI and the infosys foundations for support for the first author during his visit at HRI in December 2023, making this collaboration possible. The first author is also grateful for the excellent hospitality during his visit at HRI in December 2023. The first author was supported by Vetenskapsrådet (Swedish Science Council) within the project 2019-04890.
1. Proof of the main result
In this section we prove our main result Theorem 0.3. First we show how the condition (0.2) in [1, Theorem 3.1] can be removed, which leads to Theorem 0.2. Then we prove a lemma, which in combination with Theorem 0.2, essentially leads to Theorem 0.3.
First we have the following lemma.
Lemma 1.1.
Suppose that $\mathcal{H}$ satisfies the hypothesis in Theorem 0.3. Then there is a constant $C>0$ such that (0.2) holds.
Lemma 1.1 follows by straight-forward continuity arguments. In order to assist the reader here we present the arguments.
Proof.
Let $H_{d}=|x|^{2}-\Delta_{d}$ be the $d$-dimensional harmonic oscillator. Then $\mathscr{S}^{\prime}(\mathbf{R}^{d})$ is the inductive limit of the increasing Hilbert spaces
$\mathcal{H}_{j}\equiv H_{d}^{j}(L^{2}(\mathbf{R}^{d})),\qquad j\in\mathbf{N},$
with norms and scalar products given by
$\|f\|_{\mathcal{H}_{j}}\equiv\|H_{d}^{-j}f\|_{L^{2}(\mathbf{R}^{d})}\quad\text{and}\quad(f,g)_{\mathcal{H}_{j}}\equiv(H_{d}^{-j}f,H_{d}^{-j}g)_{L^{2}(\mathbf{R}^{d})},\quad f,g\in\mathcal{H}_{j}.$
(See e. g. [3, 13].) The fact that $\mathcal{H}$ is continuously embedded in $\mathscr{S}^{\prime}(\mathbf{R}^{d})$, implies that $\mathcal{H}$ is continuously embedded in $\mathcal{H}_{j}$, provided $j$ is chosen sufficiently large. For such $j$ we have
$|(f,\phi)|=|(f,\phi)_{L^{2}}|\lesssim\|f\|_{\mathcal{H}_{j}}\|H_{d}^{j}\phi\|_{L^{2}}\leq C\|f\|_{\mathcal{H}},$
for some constant $C>0$, and the result follows. ∎
Theorem 0.2 now follows from [1, Theorem 3.1] and Lemma 1.1.
Lemma 1.2.
[…]
  1. (1)
[…]
  2. (2)
[…]
  3. (3)
    $\|e^{i\langle\,\cdot\,,\xi\rangle}f(\,\cdot\,-x)\|=\|f\|$ for every $f\in\mathcal{H}$ and $x,\xi\in\mathbf{R}^{d}$.
Proof.
Let $Q_{R}=[-R,R]^{2d}$ be the cube in $\mathbf{R}^{2d}$ with center at origin and side length $2R$, $R\geq 1$. Then the volume of $Q_{R}$ is given by $|Q_{R}|=(2R)^{2d}$. We define the norm $\|\,\cdot\,\|_{[R]}$ by the formula
$\|f\|_{[R]}^{2}=\frac{1}{|Q_{R}|}\iint_{Q_{R}}\|e^{i\langle\,\cdot\,,\xi\rangle}f(\,\cdot\,-x)\|_{\mathcal{H}}^{2}\,dxd\xi,\qquad f\in\mathcal{H}.$
By the definition it follows that $\|\,\cdot\,\|_{[R]}$ is equivalent to $\|\,\cdot\,\|_{\mathcal{H}}$, and (0.3) gives
$C_{0}^{-1}\|f\|_{\mathcal{H}}\leq\|f\|_{[R]}\leq C_{0}\|f\|_{\mathcal{H}},\qquad f\in\mathcal{H}.$ (1.1)
Here the constant $C$ is the same as above. Furthermore, $\|\,\cdot\,\|_{[R]}$ is a norm which arises from the scalar product
$(f,g)_{[R]}=\frac{1}{|Q_{R}|}\iint_{Q_{R}}({e^{i\langle\,\cdot\,,\xi\rangle}f(\,\cdot\,-x)},{e^{i\langle\,\cdot\,,\xi\rangle}g(\,\cdot\,-x)})_{\mathcal{H}}\,dxd\xi,\quad f,g\in\mathcal{H}.$
Now let $\mho_{0}$ be a dense countable subset of $\mathcal{H}$ containing 00, and let
=$\mho=\{\,f-g,f+g,f-ig,f+ig\,;\,f,g\in\mho_{0}\,\}.$
[…]
$C_{0}^{-1}\|f_{1}\|_{\mathcal{H}}\leq\|f_{1}\|_{[m]}\leq C_{0}\|f_{1}\|_{\mathcal{H}},$
for every $m\in\mathbf{N}$, there is a subsequence $\{m_{k}\}_{k=1}^{\infty}$ of $\mathbf{N}$ such that
$\|f_{1}\|\equiv\lim_{k\to\infty}\|f_{1}\|_{[m_{k}]}$
exists. By Cantor’s diagonalization principle, there is a subsequence $\{n_{k}\}_{k=1}^{\infty}$ of $\{m_{k}\}_{k=1}^{\infty}$ such that
$\|f_{j}\|\equiv\lim_{k\to\infty}\|f_{j}\|_{[n_{k}]}$
exists for every $f_{j}\in\mho$, and by (1.1) we get
$C_{0}^{-1}\|f\|_{\mathcal{H}}\leq\|f\|\leq C_{0}\|f\|_{\mathcal{H}},$ (1.2)
when $f\in\mho$.
Next suppose that $f\in\mathcal{H}$ is arbitrary, and let $\{f_{0,j}\}_{j=1}^{\infty}\subseteq\mho_{0}$, be chosen such that
$\lim_{j\to\infty}\|f-f_{0,j}\|_{\mathcal{H}}=0.$
Then
$\|f_{0,j}-f_{0,k}\|\leq C_{0}\|f_{0,j}-f_{0,k}\|_{\mathcal{H}}\to 0,\quad\text{as}\quad j,k\to\infty.$
Since $\left|\|f_{0,j}\|-\|f_{0,k}\|\right|\leq\|f_{0,j}-f_{0,k}\|$, it follows that $\{\|f_{0,j}\|\}_{j=1}^{\infty}$ is a Cauchy sequence in $\mathbf{R}$. Hence
$\|f\|\equiv\lim_{j\to\infty}\|f_{0,j}\|$
exists and is independent of the chosen particular sequence $\{f_{0,j}\}_{j=1}^{\infty}$. Since (1.2) holds for any $f_{0,j}\in\mho$, it follows from the recent estimates and limit properties (1.2) extends to any $f\in\mathcal{H}$. This gives (1).
For $f,g\in\mathcal{H}$, their scalar product $(f,g)_{[n_{k}]}$ can be evaluated by
$(f,g)_{[n_{k}]}=\frac{1}{4}\left(\|f+g\|_{[n_{k}]}^{2}-\|f-g\|_{[n_{k}]}^{2}+i\|f+ig\|_{[n_{k}]}^{2}-i\|f-ig\|_{[n_{k}]}^{2}.\right)$
[…]
$(f,g)\equiv\frac{1}{4}\left(\|f+g\|^{2}-\|f-g\|^{2}+i\|f+ig\|^{2}-i\|f-ig\|^{2}\right),$
[…]
It remains to prove that (3) holds. By repetition it suffices to prove
$\|e^{i\langle\,\cdot\,,\xi\rangle}f(\,\cdot\,-x)\|=\|f\|$ (1.3)
when $x=0$ and $\xi=\xi_{j}e_{j}$ or when $\xi=0$ and $x=x_{j}e_{j}$ for some $x_{j},\xi_{j}\in\mathbf{R}$. Here $e_{j}$ denotes the vector of order $j$ in the standard basis of $\mathbf{R}^{d}$. Then we only prove (1.3) for $\xi=0$ and $x=x_{j}e_{j}$. The other cases follow by similar arguments and are left for the reader.
Let $R$ be chosen such that $R>|x_{j}|$. Then $Q_{R}$ and $-x_{j}e_{j}+Q_{R}$ intersects. We have
$\|f(\,\cdot\,-x_{j}e_{j})\|_{[R]}^{2}$ $=\frac{1}{|Q_{R}|}\iint_{Q_{R}}\|e^{i\langle\,\cdot\,,\eta\rangle}f(\,\cdot\,-(x_{j}e_{j}+y))\|_{\mathcal{H}}^{2}\,dyd\eta$
$=\frac{1}{|Q_{R}|}\iint_{-x_{j}e_{j}+Q_{R}}\|e^{i\langle\,\cdot\,,\eta\rangle}f(\,\cdot\,-y)\|_{\mathcal{H}}^{2}\,dyd\eta$
$=\frac{1}{|Q_{R}|}\iint_{Q_{R}}\|e^{i\langle\,\cdot\,,\eta\rangle}f(\,\cdot\,-y)\|_{\mathcal{H}}^{2}\,dyd\eta+E_{R}(f)$
$=\|f\|_{[R]}^{2}+E_{R}(f),$ (1.4)
where
$E_{R}(f)=\frac{1}{|Q_{R}|}\iint_{-x_{j}e_{j}+Q_{R}}\|e^{i\langle\,\cdot\,,\eta\rangle}f(\,\cdot\,-y)\|_{\mathcal{H}}^{2}\,dyd\eta\\ -\frac{1}{|Q_{R}|}\iint_{Q_{R}}\|e^{i\langle\,\cdot\,,\eta\rangle}f(\,\cdot\,-y)\|_{\mathcal{H}}^{2}\,dyd\eta.$
If
$\Delta_{R}=\big{(}-x_{j}e_{j}+Q_{R})\setminus Q_{R}\big{)}\bigcup\big{(}Q_{R}\setminus-x_{j}e_{j}+Q_{R}\big{)},$
then it follows that $|\Delta_{R}|=2|x_{j}|(2R)^{2d-1}$ and that
$|E_{R}(f)|\leq\frac{1}{|Q_{R}|}\iint_{\Delta_{R}}\|e^{i\langle\,\cdot\,,\eta\rangle}f(\,\cdot\,-y)\|_{\mathcal{H}}^{2}\,dyd\eta.$
This gives
$|E_{R}(f)|$ $\leq\frac{1}{|Q_{R}|}\iint_{\Delta_{R}}\|e^{i\langle\,\cdot\,,\eta\rangle}f(\,\cdot\,-y)\|_{\mathcal{H}}^{2}\,dyd\eta$
[…]
$=\frac{C_{0}|\Delta_{R}|\|f\|_{\mathcal{H}}^{2}}{|Q_{R}|}=\frac{C_{0}|x_{j}|\|f\|_{\mathcal{H}}^{2}}{R},$
which tends to zero as $R$ turns to infinity.
By letting $R=n_{k}$ in the previous analysis, (1.4) gives
$\|f(\,\cdot\,-x_{j}e_{j})\|$ $=\lim_{k\to\infty}\|f(\,\cdot\,-x_{j}e_{j})\|_{[n_{k}]}^{2}$
$=\lim_{k\to\infty}\left(\frac{1}{|Q_{n_{k}}|}\iint_{Q_{n_{k}}}\|e^{i\langle\,\cdot\,,\eta\rangle}f(\,\cdot\,-y)\|_{\mathcal{H}}^{2}\,dyd\eta+E_{n_{k}}(f)\right)$
$=\|f\|+0,$
which gives (3) and thereby the result. ∎
Proof of Theorem 0.3.
By Lemma 1.2, we may replace the norm for $\mathcal{H}$ by an equivalent norm which is invariant under translations and modulations. The result now follows from Theorem 0.2. ∎
2. Extensions and variations
In this section we slightly improve Theorem 0.3, and show that the result is still true when $\mathcal{H}$ is continuously embedded in certain distribution spaces and ultra-distribution spaces, containing $\mathscr{S}^{\prime}(\mathbf{R}^{d})$. In the first part we recall the definition of the Pilipović ultra-distribution space $\mathcal{H}_{\flat}^{\prime}(\mathbf{R}^{d})$ (also denoted by $\mathcal{H}_{\flat_{1}}^{\prime}(\mathbf{R}^{d})$ in [13]). Thereafter we extend Theorem 0.3.
We recall that the Pilipović space $\mathcal{H}_{\flat}(\mathbf{R}^{d})$ consists of all Hermite function expansions
$f(x)=\sum_{\alpha\in\mathbf{N}^{d}}c(f,\alpha)h_{\alpha}(x),$ (2.1)
where the Hermite coefficients $c(f,\alpha)$ should satisfy
$|c(f,\alpha)|\lesssim h^{|\alpha|}\alpha!^{-\frac{1}{2}},$
for some $h>0$ (see [13]). The distribution space $\mathcal{H}_{\flat}^{\prime}(\mathbf{R}^{d})$ can be identified with the set of all formal expansions in (2.1) such that
$|c(f,\alpha)|\lesssim h^{|\alpha|}\alpha!^{\frac{1}{2}},$
for every $h>0$. The distribution action between $\mathcal{H}_{\flat}(\mathbf{R}^{d})$ and $\mathcal{H}_{\flat}^{\prime}(\mathbf{R}^{d})$ is then given by
$(f,\phi)_{L^{2}}=\sum_{\alpha\in\mathbf{N}^{d}}c(f,\alpha)\overline{c(\phi,\alpha)}.$ (2.2)
We have now the following extension of Theorem 0.3.
Theorem 2.1.
Let $\mathcal{H}$ be a Hilbert space which is continuously embedded in $\mathscr{D}^{\prime}(\mathbf{R}^{d})$ or in $\mathcal{H}_{\flat}^{\prime}(\mathbf{R}^{d})$, and such that (0.3) holds true for some constant $C_{0}\geq 1$ which is independent of $f\in\mathcal{H}$ and $x,\xi\in\mathbf{R}^{d}$. If $\mathcal{H}$ is non-trivial, then $\mathcal{H}=L^{2}(\mathbf{R}^{d})$, and (0.4) holds for some constant $C>0$ which is independent of $f\in\mathcal{H}=L^{2}(\mathbf{R}^{d})$.
The proof of Theorem 2.1 follows by a combination of [15, Proposition 1.5], Theorem 0.3, and Proposition 2.2 below. First we need some preparations.
Let $\mathcal{B}$ be a vector space. A quasi-norm on $\mathcal{B}$ is a map from $\mathcal{B}$ to $\mathbf{R}_{+}\cup\{0\}$ such that
$\|\alpha f\|_{\mathcal{B}}$ $=|\alpha|\cdot\|f\|_{\mathcal{B}},$ $\qquad f$ $\in\mathcal{B},\ \alpha\in\mathbf{C}$
and
$\|f+g\|_{\mathcal{B}}^{p}$ $\leq\|f\|_{\mathcal{B}}^{p}+\|g\|_{\mathcal{B}}^{p},$ $\qquad f,g$$\in\mathcal{B}.$
If the topology on $\mathcal{B}$ is defined through this quasi-norm, then $\mathcal{B}$ is called a quasi-normed space. A complete quasi-normed space is called a quasi-Banach space.
Next we recall some facts concerning the Gelfand-Shilov space $\Sigma_{1}(\mathbf{R}^{d})$ and its distribution space $\Sigma_{1}^{\prime}(\mathbf{R}^{d})$, given in e. g. [10, 13]. We recall that $\Sigma_{1}(\mathbf{R}^{d})$ consists of all $f\in C^{\infty}(\mathbf{R}^{d})$ such that for every $h>0$, there is a constant $C_{h}>0$ such that
$|x^{\alpha}\partial^{\beta}f(x)|\leq C_{h}h^{|\alpha+\beta|}\alpha!\beta!\,.$
The smallest choice of $C_{h}$ defines a semi-norm on $\Sigma_{1}(\mathbf{R}^{d})$, and by defining a (project limit) topology from these semi-norms, it follows that $\Sigma_{1}(\mathbf{R}^{d})$ is a Fréchet space.
There are several characterizations of $\Sigma_{1}(\mathbf{R}^{d})$ and its (strong) dual or distribution space $\Sigma_{1}^{\prime}(\mathbf{R}^{d})$. For example, $\Sigma_{1}(\mathbf{R}^{d})$ is the set of all expansions in (2.1) such that
$|c(f,\alpha)|\leq C_{r}e^{-r|\alpha|^{\frac{1}{2}}}$
for every $r>0$. The smallest $C_{r}$ defines a semi-norm for $\Sigma_{1}(\mathbf{R}^{d})$, and the (project limit) topology in this setting agrees with the topology for $\Sigma_{1}(\mathbf{R}^{d})$. We may then identify $\Sigma_{1}^{\prime}(\mathbf{R}^{d})$ with all expansions in (2.1) such that
$|c(f,\alpha)|\leq Ce^{r_{0}|\alpha|^{\frac{1}{2}}},$
for some constants $C>0$ and $r_{0}>0$, with distribution action given by (2.2) when $f\in\Sigma_{1}^{\prime}(\mathbf{R}^{d})$ and $\phi\in\Sigma_{1}(\mathbf{R}^{d})$.
In the same way we may identify $\mathscr{S}(\mathbf{R}^{d})$ and $\mathscr{S}^{\prime}(\mathbf{R}^{d})$ as the sets of all expansions in (2.1) such that
$|c(f,\alpha)|\leq C_{r}(1+|\alpha|)^{-r}$
for every $r>0$, and
$|c(f,\alpha)|\leq C(1+|\alpha|)^{r_{0}}$
for some $C>0$ and $r>0$, respectively (see e. g. [11]). The distribution action is given by (2.2) when $f\in\mathscr{S}^{\prime}(\mathbf{R}^{d})$ and $\phi\in\mathscr{S}(\mathbf{R}^{d})$.
From these identifications it is evident that
$\mathcal{H}_{\flat}(\mathbf{R}^{d})\subseteq\Sigma_{1}(\mathbf{R}^{d})\subseteq\mathscr{S}(\mathbf{R}^{d})\subseteq\mathscr{S}^{\prime}(\mathbf{R}^{d})\subseteq\Sigma_{1}^{\prime}(\mathbf{R}^{d})\subseteq\mathcal{H}_{\flat}^{\prime}(\mathbf{R}^{d}),$ (2.3)
with dense and continuous embeddings.
In [13], the images of the spaces (2.3) under the Bargmann transform are presented. Let $A(\mathbf{C}^{d})$ be the set of entire functions on $\mathbf{C}^{d}$, and let
$\mathcal{A}_{0,\infty}^{\prime}(\mathbf{C}^{d})$ $\equiv\bigcup_{r\geq 0}\{\,F\in A(\mathbf{C}^{d})\,;\,|F(z)|\leq Ce^{\frac{1}{2}|z|^{2}}(1+|z|)^{r}\ \text{for some}\ C,r>0\,\}$
$\mathcal{A}_{0,1}^{\prime}(\mathbf{C}^{d})$ $\equiv\bigcup_{r\geq 0}\{\,F\in A(\mathbf{C}^{d})\,;\,|F(z)|\leq Ce^{\frac{1}{2}|z|^{2}+r|z|}\ \text{for some}\ C,r>0\,\}$
and
$\mathcal{A}_{\flat}^{\prime}(\mathbf{C}^{d})$ $\equiv A(\mathbf{C}^{d}).$
We equip $\mathcal{A}_{0,\infty}^{\prime}(\mathbf{C}^{d})$ and $\mathcal{A}_{0,1}^{\prime}(\mathbf{C}^{d})$ with inductive limit topologies through the semi-norms
$\|F\|_{\mathcal{A}_{0,\infty}^{\prime}\,,r}$ $\equiv\|F\cdot e^{-\frac{1}{2}|\,\cdot\,|^{2}}(1+|\,\cdot\,|)^{-r}\|_{L^{\infty}},$ $\quad r$ $>0,$
and
$\|F\|_{\mathcal{A}_{0,1}^{\prime}\,,r}$ $\equiv\|F\cdot e^{-(\frac{1}{2}|z|^{2}+r|\,\cdot\,|)}\|_{L^{\infty}},$ $\quad r$ $>0,$
respectively. We also equipp $\mathcal{A}_{\flat}^{\prime}(\mathbf{C}^{d})=A(\mathbf{C}^{d})$ with the projective limit topology, given by the semi-norms
$\|F\|_{\mathcal{A}_{\flat}^{\prime}\,,r}$ $\equiv\|F\|_{B_{r}(0)},$ $\quad r$ $>0.$
In [13] it is proved that the Bargmann mappings
$\mathfrak{V}_{d}$ ::: $\,\mathscr{S}^{\prime}(\mathbf{R}^{d})$ $\to\mathcal{A}_{0,\infty}^{\prime}(\mathbf{C}^{d}),$ (2.4)
$\mathfrak{V}_{d}$ ::: $\Sigma_{1}^{\prime}(\mathbf{R}^{d})$ $\to\mathcal{A}_{0,1}^{\prime}(\mathbf{C}^{d})$ (2.5)
and
$\mathfrak{V}_{d}$ ::: $\mathcal{H}_{\flat}^{\prime}(\mathbf{R}^{d})$ $\to\mathcal{A}_{\flat}^{\prime}(\mathbf{C}^{d})$ (2.6)
are homeomorphisms. Finally we recall that in terms of the Weyl map
$(W_{w}F)(z)\equiv e^{-\frac{1}{2}|w|^{2}+(z,w)}F(z-w),\quad F\in A(\mathbf{C}^{d}),\ z,w\in\mathbf{C}^{d},$ (2.7)
the Bargmann transform transfer translations and modulations as
$(\mathfrak{V}_{d}(e^{i\langle\,\cdot\,,\xi\rangle}f(\,\cdot\,-x)))(z)$ $=e^{\frac{i}{2}\langle x,\xi\rangle}(W_{\overline{w}/\sqrt{2}}(\mathfrak{V}_{d}f))(z),$ (2.8)
$z$ $\in\mathbf{C}^{d},\ w=x+i\xi\in\mathbf{C}^{d}.$
Proposition 2.2.
Let $\mathcal{B}$ be a quasi-Banach space which is continuously embedded in $\mathcal{H}_{\flat}^{\prime}(\mathbf{R}^{d})$, and such that for some constant $C>0$ and function $v\in L^{\infty}_{loc}(\mathbf{R}^{2d};\mathbf{R}_{+})$, it holds
$\|e^{i\langle\,\cdot\,,\xi\rangle}f(\,\cdot\,-x)\|_{\mathcal{B}}\leq Cv(x,\xi)\|f\|_{\mathcal{B}},\quad x,\xi\in\mathbf{R}^{d}.$ (2.9)
Then the following is true:
  1. (1)
    $\mathcal{B}$ is continuously embedded in $\Sigma_{1}^{\prime}(\mathbf{R}^{d})$;
  2. (2)
    if in addition $v(x,\xi)\leq C_{0}(1+|x|+|\xi|)^{N}$, for some constants $C_{0}>0$ and $N>0$, then $\mathcal{B}$ is continuously embedded in $\mathscr{S}^{\prime}(\mathbf{R}^{d})$.
Proof.
We may assume that $C\geq 1$. Let
$v_{0}(x,\xi)\equiv\sup_{f\in\mathcal{B}\setminus 0}\left(\frac{\|e^{i\langle\,\cdot\,,\xi\rangle}f(\,\cdot\,-x)\|_{\mathcal{B}}}{\|f\|_{\mathcal{B}}}\right).$
Then $v_{0}\leq Cv$, and $v_{0}$ is submultiplicative, i. e.,
$v_{0}(x+y,\xi+\eta)\leq v_{0}(x,\xi)v_{0}(y,\eta),\qquad x,y,\xi,\eta\in\mathbf{R}^{d}.$ (2.10)
Furthermore, (2.9) gives
$\|e^{i\langle\,\cdot\,,\xi\rangle}f(\,\cdot\,-x)\|_{\mathcal{B}}\leq v_{0}(x,\xi)\|f\|_{\mathcal{B}},\quad x,\xi\in\mathbf{R}^{d}.$
Hence we may assume that (2.9) holds with $v_{0}$ in place of $v$, and that $C=1$. We also observe that (2.10) gives
$v_{0}(x)\leq Ce^{r(|x|+|\xi|)},\quad x,\xi\in\mathbf{R}^{d},$ (2.11)
for some constants $C>0$ and $r>0$ (see [8]).
Suppose that $f\in\mathcal{B}$ and let $F=\mathfrak{V}_{d}f$. By the assumptions, (2.8) and that (2.6) is a homeomorphism, it follows that for some $C>0$ we have
$|F(z-w)e^{-\frac{1}{2}|w|^{2}+\operatorname{Re}(z,w)}|v_{0}(\sqrt{2}\overline{w})^{-1}\leq C,\qquad z\in B_{r}(0),\ w\in\mathbf{C}^{d},$
which implies
$\sup_{z\in\mathbf{C}^{d}}\left(|F(z)|e^{-\frac{1}{2}|z|^{2}}v_{0}(-\sqrt{2}\overline{z})^{-1}\right)<\infty.$ (2.12)
By combining the latter estimate with (2.11) shows that $F\in\mathcal{A}_{0,1}^{\prime}(\mathbf{C}^{d})$. The asserted embedding in (1) now follows from the homeomorphic property of (2.5).
Next suppose that $v$ satisfies the estimates in (2). Then (2.12) shows that $F\in\mathcal{A}_{0,\infty}^{\prime}(\mathbf{C}^{d})$. The asserted embedding in (2) now follows from the homeomorphic property of (2.4), giving the result. ∎
Remark 2.3.
By the definition of modulation spaces and their mapping properties under the Bargmann transform, it follows from the estimate (2.12) that $\mathcal{B}$ in Proposition 2.2 (1) is continuously embedded in the modulation space $M^{\infty}_{(\omega)}(\mathbf{R}^{d})$, with $\omega(x,\xi)=v_{0}(-x-i\xi)^{-1}$, $x,\xi\in\mathbf{R}^{d}$. (See e. g. [13].)
Proof of Theorem 2.1.
If $\mathcal{H}$ is continuously embedded in $\mathscr{D}^{\prime}(\mathbf{R}^{d})$ or in $\mathcal{H}_{\flat}(\mathbf{R}^{d})$, then [15, Proposition 1.5] and Proposition 2.2 (2) show that $\mathcal{H}$ is continuously embedded in $\mathscr{S}^{\prime}(\mathbf{R}^{d})$. The result now follows from Theorem 0.3. ∎